OpenStudy (anonymous):

step-by-step how do i solve 1. (x + 5)(x - 2) = 0 2. x^2 - 3x + 27

OpenStudy (alfie):

\[\large \begin{array}{l} (x + 5)(x - 2) = 0\\ {x^2} - 10 - 2x + 5x = 0\\ {x^2} + 3x - 10 = 0\\ \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\\ \frac{{ - 3 \pm \sqrt {9 + 40} }}{2}\\ \frac{{ - 3 \pm 7}}{2}\\ {x_1} = - 5,{x_2} = 2 \end{array}\]

OpenStudy (alfie):

\[\large \begin{array}{l} {x^2} - 3x + 27\\ \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\\ \frac{{3 \pm \sqrt {9 - 108} }}{2}\\ It - is - negative - so - no - solutions - in - R \end{array}\]

OpenStudy (alfie):

The root is negative, obviously :)